- ~7/31: Trial for DRMI with polarization.
- 7/6~: Trial for mass lock
- Long term:
- 7/1-8/9: FPMI/DRFPMI commissioning
- 8/12-10/4: ASC
- 10/7-11/15: Post Commissioning
- 11/18-12/13: Noise hunting
Report of work
- Temperature drift affected IFO commissioning. So tried ALS signal handling off to TRX and TRY.
- SR3 H2 noise investigation is still going on.
- Suppressing 80mHz of yaw by feeding back the MN oplev signal was done.
- OMC work for alignment. Control system was implemented.
- temperature fluctuation at all the central area specially around BS for 0.6C with 20hours period.
- DGS maintenance
- BIO cards installation.
- Replacing RT-PCs of SR or PR
- Modification of Type A model.
- Temperature maintenance in the afternoon.
- MZ work (Guiguo).
- OMC work in the morning.